Costs per Equivalent Unit and Production Costs The following information concerns production in the Forging Department...

  1. Costs per Equivalent Unit and Production Costs

    The following information concerns production in the Forging Department for November. All direct materials are placed into the process at the beginning of production, and conversion costs are incurred evenly throughout the process. The beginning inventory consists of $9,000 of direct materials.

    ACCOUNT Work in Process—Forging Department ACCOUNT NO.
    Date Item Debit Credit Balance
    Debit Credit
    Nov. 1 Bal., 900 units, 60% completed 10,566
    30 Direct materials, 12,900 units 123,840 134,406
    30 Direct labor 21,650 156,056
    30 Factory overhead 16,870 172,926
    30 Goods transferred, ? units ? ?
    30 Bal., 1,400 units, 70% completed ?

    Cost per equivalent units of $9.60 for Direct Materials and $3.00 for Conversion Costs.

    Based on the above data, determine each of the following amounts.

    If required, round your interim calculations to two decimal places. Round final answers (a-c) to the nearest dollar.

    a. Cost of beginning work in process inventory completed in November.
    $

    b. Cost of units transferred to the next department during November.
    $

    c. Cost of ending work in process inventory on November 30.
    $

    d. Costs per equivalent unit of direct materials and conversion included in the November 1 beginning work in process. If required, round your answers to two decimal places.

    Direct materials cost per equivalent unit $
    Conversion cost per equivalent unit $

    e. The November increase or decrease in costs per equivalent unit for direct materials and conversion from the previous month. If required, round your answers to two decimal places.

    Increase or Decrease Amount
    Change in direct materials cost per equivalent unit
    • Decrease
    • Increase
    $
    Change in conversion cost per equivalent unit
    • Decrease
    • Increase
    $

    Feedback

Loading item

There was an error loading this item. If this continues to occur, please contact Technical Support.
Answer #1

a. Cost of beginning work in process inventory completed in November = 11646

b. Cost of units transferred to the next department during November = 144,900 +11646 = 156,546

c. Cost of ending work in process inventory on November 30 = $16380

d.

I. Direct materials cost per equivalent unit = 9000/900 = $10 per unit

Conversion cost = (10,566-9000) / (900 *60%) = $2.9 per unit

e. Change in direct materials cost per equivalent unit = $0.4 , decrease ( $10 - 9.6)

Change in conversion cost per equivalent unit = $0.1, decrease ( 3- 2.9)

Workings:

Physical units
Beginning Units 900
Units started 12,900
Total units 13,800
Units Completed 12,400
Ending units 1,400
Total 13,800
Physical Units Direct Materials Conversion
Units from beginning WIP 900 0 360 [900*40%]
Units Completed 11500 11500 11500
Ending units 1400 1400 980 [1400*70%]
Total units accounted for 12900
Total equivalent units 12900 12840
Direct Materials Conversion Total
Current month 123,840 38,520
Total 123,840 38,520
Equivalent units 12900 12840
Equivalent cost per unit 9.6 3.0 13
Equivalent units Cost per unit Total cost
Beginning cost 10566
current period 360 3 1080
Total cost of beginning WIP 11646
Equivalent units Cost per unit Total cost
Materials 11500 9.6 110400
Conversion 11500 3 34500
Total cost of units completed this month 144900
Equivalent units Cost per unit Total cost
Materials 1400 9.6 13440
Conversion 980 3 2940
cost of ending WIP 16380